OpenStudy (anonymous):

A polynomial 2z^5+az^4+bz^3+cz^2+dz+e with real coefficients a,b,d,e has 2i, -i and 1 as roots. Find c.

ganeshie8 (ganeshie8):

\[2z^5+az^4+bz^3+cz^2+dz+e = 2(z-1)(z+2i)(z-2i)(z+i)(z-i)\]

ganeshie8 (ganeshie8):

compare z^2 coefficient both sides

ganeshie8 (ganeshie8):

maybe simplify the right hand side a bit first using : (a+bi)(a-bi) = a^2 + b^2

ganeshie8 (ganeshie8):

not sure if there is any other easy way..

OpenStudy (anonymous):

Lol, I hope to God there is xD

ganeshie8 (ganeshie8):

hmm there could be, but i feel its not worth trying for other ways as they're not so obvious lol..

ganeshie8 (ganeshie8):

i would just expand and compare coefficients both sides : \[2z^5+az^4+bz^3+cz^2+dz+e = 2(z-1)(z+2i)(z-2i)(z+i)(z-i) \] \[2z^5+az^4+bz^3+cz^2+dz+e = 2(z-1)(z^2+4)(z^2+1)\]

OpenStudy (anonymous):

Ooooooh, I see. So just open it up and see what has z^2.

OpenStudy (anonymous):

That's not too bad.

ganeshie8 (ganeshie8):

exactly ! its trivial to solve, but agree its a dumb method lol

OpenStudy (anonymous):

Thanks x)

OpenStudy (anonymous):

One question. Why z+i and z-i?

ganeshie8 (ganeshie8):

thats a very good question :) cuz of a theorem : if a is `k` root of `P(x)`, then `(x-k)` is a factor of `P(x)`

ganeshie8 (ganeshie8):

and one more theorem : complex roots always come in conjugate pairs (if `a + bi` is a root, then `a - bi` will also be a root)

OpenStudy (anonymous):

But the roots are, 2i, -i, 1. Wouldn't it be: (z-2i) ( z+i) (z-1)?

OpenStudy (anonymous):

Oh, I see.

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

That helps a lot. Gonna be a problem on my final tomorrow. Have a good weekend :)

ganeshie8 (ganeshie8):

if you're feeling lazy, ur z^2 coefficient is here : http://www.wolframalpha.com/input/?i=simplify+2%28z-1%29%28z%2B2i%29%28z-2i%29%28z%2Bi%29%28z-i%29

OpenStudy (anonymous):

Hahaha, no wolfram on exams x)

ganeshie8 (ganeshie8):

wish u all the best wid the exam :)

OpenStudy (anonymous):

thanks :)